Carla is 5 years younger than cathy,2 years ago the sum of their age is 35,how old is carla now?
Answer by phillywily(18)

Carla=x
Cathy=x+5
(x-2)+(x+5-2)=35
x-2+x+5-2=35
2x+5-4=35
2x=35-5+4
2x=34
x=17